OpenStudy (anonymous):

Find the thirtieth term of the following sequence. -6, -4, -2, 0, ... 52 64 66

OpenStudy (anonymous):

First, you must find the number series that those numbers follow. Do you know how to find this?

OpenStudy (anonymous):

Nope, im jus tryina do my work soo i can go home nd have a good B-Day

OpenStudy (anonymous):

whats the difference between a -6 and a -4? difference between -4 and -2 ? difference between -2 and 0 ?

OpenStudy (anonymous):

+2

OpenStudy (anonymous):

Happy birthday! I'll make it a simple version then. To find the number series expression, you must find the difference between each number, which in this case is 2. Now the beginning value is -6, so (something)+2n (n being the number of the term in the series. To find what the (something) is, you plug 1 in for n and solve, setting it equal to the first term: x+2(1)=-6 x=-8 So your beginning number of the number series is -8, and the difference between each number is +2, so the number series looks like this: x=-8+2n.

OpenStudy (anonymous):

the formula is : an = a1 + (n - 1) * d a1 is first term = -6 n = the term you want to find = 13 d = the difference between the numbers = 2 now plug all your stuff into the formula

OpenStudy (anonymous):

Soo thebanswer is?....., im guessing 64?

OpenStudy (anonymous):

the answer**

OpenStudy (anonymous):

x=-8+2n. Plug 30 in for n and solve.

OpenStudy (anonymous):

Okk & Ty

OpenStudy (anonymous):

a13 = -6 + (13 - 1) * 2 a13 = -6 + 12 * 2 a13 = -6 + 24 a13 = 18 <-- 13th term are you looking for the 13 term or the sum ?

OpenStudy (anonymous):

It's the 30th term.

OpenStudy (anonymous):

oh...my bad.....I feel stupid now

OpenStudy (anonymous):

Don't worry. I read it as 13th too at the beginning...

OpenStudy (anonymous):

Happy Birthday @LaDarriusWatson

OpenStudy (anonymous):

Ty ;)

OpenStudy (anonymous):

Im jus turning 16

OpenStudy (anonymous):

sorry about the mess up I made on your question......16th birthday....hmmm...don't party too hard ...lol

OpenStudy (anonymous):

Lmao